# Pushing the Envelope of LLM Inference on AI-PC

### v1

@@ III-E Performance modeling
In this section we devise roofline models for the 2-bit and 1-bit GEMM microkernels to assess their efficacy and limitations. These models will be used to explain the obtained performance results in section IV. Specifically, we aim to determine under which conditions our microkernels can saturate the available bandwidth on the platform and effectively accelerate matrix-vector multiplications during inference. For our analysis we follow the reasoning/bottleneck-style analysis of the original roofline model [25]. Let denote the number of bytes/cycle each core in a multi-core CPU can read when all cores access memory simultaneously. Let represent the number of cycles to compute an output 256-bit vector in the GEMM without considering the data load instructions. Assuming bytes must be read from memory in order to compute an output 256-bit vector, the number of cycles to calculate such an output 256-bit vector is:
@@ IV-A Experimental Platforms
Intel Core Ultra 7 258V CPU (henceforth referred to as LNL) with 8 cores (4 performance and 4 efficiency cores). It has memory DDR5@8533 MT/s with measured read bandwidth of 97 GB/s and Maximum Turbo Power of 37W.
IV-B GEMV microbenchmarks
@@
In Figure 5 we illustrate the GEMV micro-benchmark results on ARL. As a baseline we consider the int8 GEMM performance where essentially there is no upconvert overhead, and the GEMV microkernel merely consists of vector loads and vnni-int8 FMAs. The int2 GEMV kernels are within 10-20% of the int8 roofline and achieve bandwidth close to the machine’s peak. Especially for larger matrix sizes (right side of the plot), where the OpenMP and barrier overheads are less significant, the int2 GEMV achieves bandwidth within 2-5% of the int8 GEMV. For the smallest matrix shape , given the peak read bandwidth (98 Gb/s), the ideal pure int2 microkernel time is 6 microseconds, while we measure 8.9 microseconds. However, OpenMP parallel region fork/join and barrier overheads account for 1 microsecond as measured via benchmarking on ARL. In contrast, the int8 GEMV kernels move more data than the int2 ones, making the relative OpenMP overhead less impactful. These overheads are even more pronounced for 1-bit weights (dark green bars).
@@
IV-C End to end inference results
@@
For the 1B model, in the 16-bit run on ARL we get (i.e. 87% of the time is spent in the GEMVs), and as a result the maximum attainable speedup is 4.3 for the 2-bit inference () and 5.5 for the 1-bit inference (). Under this lens, we indeed see that out obtained results are close to the ideal ones (we observed 4.1 and 4.8 respectively). On the other extreme with the 8B model, in the 16-bit run we get on ARL (i.e. 96% of the time is spent in the GEMVs), and consequently the maximum attainable speedup is 6.1 for the 2-bit inference () and 9.6 for the 1-bit inference (). Again we see that our obtained results are close to the ideal ones (we observed 5.8 and 8.3 speedup over the 16-bit baseline).
@@
IV-D Comparison with 2-bit inference on GPU
@@
In a recent update, bitnet.cpp included CUDA optimized 2-bit kernels tailored for the NVIDIA A100 platform, and we were able to reproduce the reported results in that repo 111https://github.com/microsoft/BitNet/tree/main/gpu on our local A100 GPU. To have an apples-to-apples comparison with our CPU results, we benchmarked our 2-bit inference pipeline on a same-size 2 Billion parameter model222https://huggingface.co/andrijdavid/Llama3-2B-Base. On the A100 GPU and with the bitnet-b1.58 2 Billion model we were able to obtain 250 tokens/s, whereas we achieve on the aforementioned same-sized 2B model 110 tokens/s on ARL, 82 tokens/s on ARLH, and 88 tokens/s on LNL. As a result, we conclude that this level of performance on the CPUs is within 2.3-3 of the A100 GPU which has 17-20 more bandwidth than our CPUs. To further shed light on these results, we assessed the efficacy of the 2-bit GEMV kernels on A100 compared to the corresponding 16-bit ones (these results are also provided in the bitnet.cpp and we were able to reproduce them). On the A100 GPU, for the smaller matrix shapes (e.g. 25602560, 38402560, 32003200) the observed 2-bit GEMV speedup over the 16-bit GEMV is only 1.3-1.4 whereas for the larger shapes the speedup goes up to 3.6. Considering that the majority of matrices in such LLMs (i.e. 2B) are relatively small, we expect the upside of the 2-bit GEMVs to be limited on A100. For comparison, in similarly-sized matrices we were able to observe close to ideal bandwidth with our 2-bit GEMV kernels on CPUs, meaning that they are faster than the corresponding 16-bit GEMVs, and as a result we materialized substantial speedup in the end-to-end 2-bit inference.
@@ V Conclusions And Future Work
In this work, we designed and implemented 1-bit and 2-bit GEMM microkernels optimized for modern x86 CPUs, achieving close to roof-line efficiency across a variety of CPU platforms. We integrated these microkernels into a state-of-the-art LLM inference framework and presented end-to-end inference results with 2-bit models that outperform the current SOTA runtime (bitnet.cpp) by up to 2.2, and deliver up to 7 speedup compared to the 16-bit model inference. Our optimized runtime advances the state of LLM inference on AI PCs and edge devices, paving the way for efficient deployment of ultra-low-bit LLM models. Additionally we demonstrate that with proper microkernel design and runtime support, ultra-low-bit inference on CPUs can approach GPU-level performance. As future work we plan to extend our work to ARM platforms, since essentially our 1-bit and 2-bit microkernels can be readily translated to AArch64 and SVE instructions which are available on modern ARM CPUs.

### v2

@@ III-F Performance modeling & limitations of the CPU kernels
The low-precision GEMMs on CPU require additional work to up-convert the low-precision operand to int8 since there is no native hardware support e.g. for int2int8 FMA operations (unlike Xe2 GPUs). Also, the aggregate int8 FMA throughput on modern AI-PC with limited number of cores poses a challenge in achieving close to roof-line performance for low-precision CPU GEMMs. In this section we devise roofline models for the 2-bit and 1-bit CPU GEMM microkernels to assess their efficacy and limitations. These models will be used to explain the obtained performance results in section IV. Specifically, we aim to determine under which conditions our CPU microkernels can saturate the available bandwidth on the platform and effectively accelerate matrix-vector multiplications during inference. For our analysis we follow the reasoning/bottleneck-style analysis of the original roofline model [28]. Let denote the number of bytes/cycle each core in a multi-core CPU can read when all cores access memory simultaneously. Let represent the number of cycles to compute an output 256-bit vector in the GEMM without considering the data load instructions. Assuming bytes must be read from memory in order to compute an output 256-bit vector, the number of cycles to calculate such an output 256-bit vector is:
@@ IV-A Experimental Platforms
Intel Core Ultra 7 258V CPU (henceforth referred to as LNL) with 8 cores (4 performance and 4 efficiency cores). It has 16GB of DDR5@8533 MT/s memory with measured read bandwidth of 97 GB/s and Maximum Turbo Power of 37W.
For our GPU evaluation we use two Xe2 GPU systems:
Intel Arc 140V, which is the integrated GPU in the aforementioned Lunar Lake (LNL) platform. It consists of 8 Xe2 GPU cores and it has access to the same memory/bandwidth as the LNL CPU (16 GB of DDR5@8533 MT/s with measured read bandwidth of 97 GB/s).
Intel Battlemage Arc B580 (BMG) discrete GPU. It consists of 20 Xe2 cores, and it has 12 GB of GDDR6 memory, delivering up to 456 GB/s. The int8 peak is 233 TOPS (Trillions Operations per Second) and its Total Board Power (TBP) is 190W.
IV-B GEMV microbenchmarks on CPU
@@
In Figure 5 we illustrate the GEMV micro-benchmark results on ARL. As a baseline we consider the int8 GEMM performance where essentially there is no upconvert overhead, and the GEMV microkernel merely consists of vector loads and vnni-int8 FMAs. The int2 GEMV kernels are within 10-20% of the int8 roofline and achieve bandwidth close to the machine’s peak. Especially for larger matrix sizes (right side of the plot), where the OpenMP and barrier overheads are less significant, the int2 GEMV achieves bandwidth within 2-5% of the int8 GEMV. For the smallest matrix shape , given the peak read bandwidth (98 Gb/s), the ideal pure int2 microkernel time is 6 microseconds, while we measure 8.9 microseconds. However, OpenMP parallel region fork/join and barrier overheads account for 1 microsecond as measured via benchmarking on ARL. In contrast, the int8 GEMV kernels move more data than the int2 ones, making the relative OpenMP overhead less important. These overheads are even more pronounced for 1-bit weights (dark green bars).
@@
IV-C GEMV/GEMM microbenchmarks on Intel Xe2 GPUs
In this section we present analogous GEMV microbenchmarks for our Xe2 GPU systems. In addition to the memory bound GEMV microbenchmarks (), we present here microbenchmarks results for GEMM shapes with larger dimension that should lead to compute bound kernels. The contemporary Xe2 systems offer high compute throughput via the int2int8int32 DPAS instructions (same as the int8int8int32 throughput, e.g. TFLOPS for the integrated Intel Arc 140V and TFLOPS for Battlemage B580), and as such should be able to accelerate the compute bound/prefill phases of the int2 inference compared to the pure BF16 inference. Therefore, we will also assess the efficacy of our mixed precision int2 BF16BF16 Xe2 GPU kernels for compute-bound GEMM cases.
In Figure 8 we illustrate the GEMV micro-benchmark results on the integrated Intel Arc 140V (LNL-GPU). As a baseline (blue bars) we consider the int8int8 GEMV performance from oneDNN where essentially there is no input/output quantization/dequantization overheads. The int2 BF16BF16 GEMV kernels (including the calculation of the scales) are within 10% of the int8 roofline and achieve bandwidth close to the machine’s peak (see orange bars). Especially for larger matrix sizes (right side of the plot), where the kernel launch overheads are less significant, the int2 GEMV achieves bandwidth same as the int8 GEMV. For the smallest matrix shape , given the peak read bandwidth (98 Gb/s), the ideal pure int2 microkernel time is 6 microseconds. However, the kernel launch overhead is in the order of 2 microseconds, thus we effectively observe 53 Gb/s bandwidth. We also observe that the sustained bandwidth by the int2 Xe2 LNL-GPU kernels is comparable and slightly higher than the to the one achieved by the LNL-CPU cores (see Figure 7). This is expected since both CPU and Xe2 CPU have access to the same DDR5 memory and the Xe2 kernels with the native int2int8 DPAS instructions are able to alleviate some of the upconvert overheads that exist on the CPU kernels. We note that for all the GEMV cases, we found experimentally that fusing the scale calculation within the kernel (in order to avoid separate kernel launches at the expense of redundant computation) is beneficial, especially since the tensor is merely an activation vector. All these results indicate that our fused BF16BF16 Xe2 kernels described in Section III-E are able to effectively hide the quantization/dequantization overheads an achieve close to peak memory bandwidth.
Figure 9 shows the attained GEMM performance () for the same weight matrices as the ones in the previous microbenchmarks. These GEMM have substantially higher weight re-use compared to the GEMV cases and in principle should be compute bound. To shed light on the obtained performance we illustrate the obtained performance in TFLOPS for 3 different configurations: i) a pure int8int8 GEMM from oneDNN without any quantization/dequantization overheads that should serve as attainable roofline (blue bars), ii) the fused int2 BF16BF16 Xe2 GEMM kernel only (illustrated in Figure 3, i.e. without the kernel/pass for calculating the scales) depicted with orange bars, and iii) the full-fledged mixed precision int2 GEMM including the overhead of calculating the scales via a separate kernel (green bars). First of all, we observe that the fused mixed precision int2 GPU GEMM kernels described in Section III-E achieve performance close to the pure int8int8 GEMM (note that the int2 int8 DPAS instruction has the same throughput as the int8 int8 DPAS instruction). This result highlights the efficacy of our fused implementation that essentially hides all the online quantization and dequantization computations. For these compute-bound GEMMs where the input activations are not simply a vector, we found experimentally that the best performance is obtained by performing the input scale calculation as a separate kernel before the GEMM (instead of fusing this scale calculation in the GEMM kernel like we did in the GEMV cases). This overhead is typically in the range of 5-15% and for the cases with larger it is diminished down to 3%. Finally comparing the full mixed precision kernel performance (green bars) with the performance of the pure int8 kernels from oneDNN serving as an attainable roofline (blue bars), we observe that the int2 BF16BF16 Xe2 GEMMs are within 5-15% of the pure int8 oneDNN GEMMs. For the larger tested matrices/shapes, we achieve on LNL-GPU up to 47 TFLOPS with the full-fledged int2 GEMM (whereas the maximum achieved is 53 TFLOPS by the int8 GEMM, essentially being close to 90% of the attainable peak). For comparison, LNL-CPU cores have a peak aggregate int8 FMA throughput of TFLOPS, therefore using the int2 GEMM kernels on LNL-GPU Xe2 cores enable substantial benefits for compute bound use-cases (e.g. prefill inference stage, larger batch size).
Figures 10 and 11 illustrate the analogous GEMV and GEMM results on the BMG discrete GPU, and the conclusions are similar to the ones for the LNL-GPU case: For both GEMV and GEMM cases the mixed precision int2 BF16BF16 Xe2 GEMM achieves performance close to the pure int8int8 GEMM from oneDNN which serves as an attainable roofline. Especially for larger matrices (right side of the plots), the obtained performance for the ultra-low bit kernels is within 5-10% of the roofline, effectively showing that our fused kernel implementation diminishes the necessitated quantization and dequantization overheads. For the larger GEMV cases occurring in Llama3-8B model, BMG achieves bandwidth up to 380 GB/s and for the corresponding GEMM cases () we achieve up to 175 TFLOPS via our mixed precision implementation. We note here that due to the increased bandwidth of BMG (450 Gb/s vs 97 GB/s on LNL-GPU) the kernel launch overheads are more emphasized for the cases where the involved matrices are small (e.g. MobileLLM-1.5B model).
IV-D End to end inference results on CPU
@@
For the 1B model, in the 16-bit run on ARL we get (i.e. 87% of the time is spent in the GEMVs), and as a result the maximum attainable speedup is 4.3 for the 2-bit inference () and 5.5 for the 1-bit inference (). Under this lens, we indeed see that our obtained results are close to the ideal ones (we observed 4.1 and 4.8 respectively). On the other extreme with the 8B model, in the 16-bit run we get on ARL (i.e. 96% of the time is spent in the GEMVs), and consequently the maximum attainable speedup is 6.1 for the 2-bit inference () and 9.6 for the 1-bit inference (). Again we see that our obtained results are close to the ideal ones (we observed 5.8 and 8.3 speedup over the 16-bit baseline).
@@
IV-E End to end inference results on Intel Xe2 GPUs
For the evaluation of the end to end inference with the Intel Xe2 GPUs, we integrated our optimized kernels into the vLLM [5] framework as a quantization plugin and evaluated the inference on the same datasets and setup described in Section IV-D. Figure 15 exhibits the end to end inference on LNL-GPU with BF16 precision (blue bars) and 2-bit weights (green bars), and we observe that the 2-bit inference provides speedup up to 6.3 for the inference with the Llama3-8B model, achieving for this case 29.2 tokens/second. As expected, this throughput is in alignment with the obtained LNL-CPU result (28 tokens/sec) since both CPU and Xe2 GPU on the LNL platform have access to the same bandwidth. However, on the LNL AI-PC platform one can benefit from leveraging the Xe2 GPU in compute-bound use-cases like inference with longer context, larger batch size and shorter output length, since the compute-bound GEMM performance gap of LNL-CPU and LNL-GPU is 30 as illustrated in Section IV-C. We ran an experiment with the 2-bit Llama3-8B model, 2048 input tokens and 16 output tokens, and we observed that LNL-GPU is faster than LNL-CPU with respect to the end to end latency.
Figure 16 illustrates the end to end inference on BMG for the same experimental setup. The largest BF16 model (Llama3-8B) does not fit in the memory of BMG (12 GB) and as such we denote that point with “OOM” (Out-Of-Memory) in the plot. On the other hand, our 2-bit Llama3-8B model requires only 2 GB of memory, and on BMG the int2 end to end inference with our Xe2 kernels obtains 135 tokens/sec. As a sanity check, this performance on BMG is 4.6 higher than the one obtained on LNL-GPU, which is in alignment with their relative 4.7 gap in the bandwidth. For the case with the smaller MobileLLM-1.5B model, the observed speedup on BMG is only 1.1. We shed light on this behavior by looking at the detailed breakdown of the inference. For this model, our detailed breakdown shows that the GEMV kernels on the GPU device are indeed accelerated by 3 (compared to BF16) which is in alignment with the GEMV microbenchmarks of Figure 10 (middle bars). However, end-to-end inference latency has one more component, where CPU scheduling and synchronization time plays a critical role. Our optimized int2 Xe2 kernels reduce the GPU device time substantially, and as a result the CPU scheduling time gets exposed, impacting the overall speedup. More specifically, the CPU overheads for such cases with very small involved matrices are increased by 3.6 compared to the BF16 inference. These CPU overheads are exposed and can not be hidden fully behind the int2 GEMV kernels with very small device time, negating the benefits obtained in pure device time in GEMV.
IV-F CPU comparison with 2-bit inference on NVIDIA GPU
@@
In a recent update, bitnet.cpp included CUDA optimized 2-bit kernels tailored for the NVIDIA A100 platform, and we were able to reproduce the reported results in that repo111https://github.com/microsoft/BitNet/tree/main/gpu on our local A100 GPU. To have an apples-to-apples comparison with our CPU results, we benchmarked our 2-bit inference pipeline on a same-size 2 Billion parameter model222https://huggingface.co/andrijdavid/Llama3-2B-Base. On the A100 GPU and with the bitnet-b1.58 2 Billion model we were able to obtain 250 tokens/s, whereas we achieve on the aforementioned same-sized 2B model 110 tokens/s on ARL, 82 tokens/s on ARLH, and 88 tokens/s on LNL. As a result, we conclude that this level of performance on the CPUs is within 2.3-3 of the A100 GPU which has 17-20 more bandwidth than our CPUs. To further shed light on these results, we assessed the efficacy of the 2-bit GEMV kernels on A100 compared to the corresponding 16-bit ones (these results are also provided in the bitnet.cpp and we were able to reproduce them). On the A100 GPU, for the smaller matrix shapes (e.g. 25602560, 38402560, 32003200) the observed 2-bit GEMV speedup over the 16-bit GEMV is only 1.3-1.4 whereas for the larger shapes the speedup goes up to 3.6 (see Figure 17, bf16 GEMV performance on A100 with yellow bars and int2 GEMV on A100 with green bars). Considering that the majority of matrices in such LLMs (i.e. 2B) are relatively small, we expect the upside of the 2-bit GEMVs to be limited on A100. For comparison, in similarly-sized matrices we were able to observe close to ideal bandwidth with our 2-bit GEMV kernels on CPUs, meaning that they are faster than the corresponding 16-bit GEMVs, and as a result we materialized substantial speedup in the end-to-end 2-bit inference.
IV-G BMG comparison with 2-bit inference on NVIDIA GPU
Finally, we compare the 2-bit end to end inference performance on B580 (BMG) with the A100 2-bit inference from bitnet.cpp on a same-sized 2 B model (same setup as the one in Section IV-F). On BMG we obtained 365 tokens/sec, essentially being 1.46 faster than the 2-bit inference on A100 despite the fact that BMG has almost 4 less bandwidth than A100. To understand this behavior, we illustrate in Figure 17 the GEMV execution times for various shapes and precisions on both BMG and A100. First by comparing the BF16 GEMV times (magenta bars for BMG and yellow bars for A100) we observe that A100 is for larger matrices 3-4 faster than BMG, which aligns with the 4 gap in the bandwidth among the two platforms (see green dotted speedup line). However, when comparing the int2/2-bit GEMV times (blue bars for BMG and green bars for A100) the landscape changes: BMG is up to 2.2 faster than A100 (see blue dotted speedup line) despite having less available bandwidth. The efficiency of our 2-bit Xe2 GEMV kernels (e.g. see Figure 10) is substantially better than the ones of the 2-bit A100 GEMV kernels (which show only limited speedup of 1.3-1.4 over the BF16 GEMV for smaller matrix shapes). As a result, the 2-bit end-to-end inference on BMG ends up being faster than the 2-bit inference on A100 for a 2B parameter model.
@@ V Conclusions And Future Work
In this work, we designed and implemented 1-bit and 2-bit GEMM microkernels optimized for modern x86 CPUs, achieving close to roof-line efficiency across a variety of CPU platforms. We integrated these microkernels into a state-of-the-art LLM inference framework and presented end-to-end inference results with 2-bit models that outperform the current SOTA runtime (bitnet.cpp) by up to 2.2, and deliver up to 7 speedup compared to the 16-bit model inference. We then extend this work to Intel GPUs where we design and implement mixed precision, 2-bit GEMM kernels, and show their performance to be close to optimal, speeding up the BF16 inference up to 6.3. On a discrete Intel Arc B580 Xe2 GPU we achieve a speedup of 1.5 over the 2-bit inference on A100 for a same-sized 2 Billion parameter model, despite A100 having 4 more bandwidth than B580. Thus, our work pushes the envelope of LLM inference on AI-PCs and Intel Xe GPUs, and demonstrates that ultra-low-bit inference even on CPUs and discrete client GPUs can approach high-end GPU-level performance. As future work we plan to extend our work to ARM platforms, since essentially our 1-bit and 2-bit microkernels can be readily translated to AArch64 and SVE instructions which are available on modern ARM CPUs.